# Learning Laplacian Eigenvectors: a Pre-training Method for Graph Neural Networks

### v1

@@ Abstract
We propose a novel framework for pre-training Graph Neural Networks (GNNs) by inductively learning Laplacian eigenvectors. Traditional Message Passing Neural Networks (MPNNs) often struggle to capture global and regional graph structure due to over-smoothing risk as network depth increases. Because the low-frequency eigenvectors of the graph Laplacian matrix encode global information, pre-training GNNs to predict these eigenvectors encourages the network to naturally learn large-scale structural patterns over each graph. Empirically, we show that models pre-trained via our framework outperform baseline models on a variety of graph structure-based tasks. While most existing pre-training methods focus on domain-specific tasks like node or edge feature reconstruction, our self-supervised pre-training framework is structure-based and highly flexible. Eigenvector-learning can be applied to all graph-based datasets, and can be used with synthetic features when task-specific data is sparse.
@@ 1 Introduction
Graph Neural Networks (GNNs) have become a powerful tool in node and graph representation learning, with successful applications across domains ranging from biomedicine [1, 2, 3, 4] to social networks [5]. However, traditional message passing GNNs often struggle to capture the global and regional structure of the graph, primarily due to the risk of over-smoothing as network depth increases [6, 7]. One promising solution is to pre-train GNNs on structure-aware tasks [4]. In particular, pre-training a GNN to produce the low-frequency eigenvectors of the graph Laplacian enables it to learn global patterns within the data. These low-frequency eigenvectors capture global and regional structure; by training the GNN to produce the lower frequency components, we encourage it to focus on the larger-scale structure of the graph. The pre-trained model can then be fine-tuned on other downstream, context-aware tasks, or its embeddings can be used as positional encodings for Graph Transformer networks. Moreover, since Laplacian eigenvector learning is entirely structure-based, it can be used as a training task for transfer learning and graph foundation models.
While some structure-informed pre-training tasks have been proposed, the majority are tailored to specific domains and rely on access to informative node features [3, 2]. Common strategies include context-prediction and motif or cluster prediction [3, 4, 2]. Existing domain-agnostic pre-training strategies also suffer from over-smoothing, as they rely solely on deep message passing GNNs [3, 1, 2] with basic node-wise prediction heads. This highlights the need for an alternative prediction head that can naturally capture and synthesize long-range structural information.
@@
    We introduce Laplacian eigenvector-learning as a structure-based pre-training task for GNNs.
@@
    We introduce a global MLP head that enables long-range interaction between vertices within the graph, improving the effectiveness of eigenvector-based pre-training.
@@
    We demonstrate that our pre-training framework provides performance improvements over baseline models on various structure-based downstream applications.
2 Background and Notation
The Laplacian $L$ of a graph $G$ is defined as:
L=DAL=D-A
where $D$ is the diagonal degree matrix and $A$ is the unnormalized adjacency matrix of $G$.
Let $\lambda_{1},\lambda_{2},\dots\lambda_{k}$ denote the $k$ lowest eigenvalues of $L$ in nondecreasing order. Let $\psi_{1},\psi_{2},\dots\psi_{k}$ denote the corresponding eigenvectors.
By Courant-Fischer, the eigenvectors of $L$ (and the eigenvectors of any Hermitian matrix) can be expressed as the following iterative optimization problem:
$\psi_{k}\in\operatorname*{argmin}_{\begin{subarray}{c}\|x\|=1\\ x\perp\psi_{1},\dots,\psi_{k-1}\end{subarray}}x^{\top}Lx.$
@@ 3 Method
3.1 Model Architecture
@@
Graph-level MLP: We concatenate the node-wise output of the base GNN model to form a graph-level aggregated representation. We then pass the aggregated vector through an MLP model to produce the low-frequency Laplacian eigenvectors. Concatenating the node embeddings prior to applying the MLP allows the model to learn relationships between distant nodes without risking over-smoothing.
Previous eigenvector-learning methods use a node-wise MLP head, processing each node’s eigencoordinates independently based on their learned hidden embedding [8, 9, 1].
3.2 Node Feature Augmentation
To provide the model with additional structural information, we augment node features with structure-based embeddings. We propose two kinds of embeddings: (1) wavelet positional embeddings, which encode relative positional information between nodes, and (2) diffused dirac embeddings, which encode local connectivity structures around each node. Both embeddings use the random walk matrix, and capture local aggregate information on each node. We discuss these embeddings in detail in A.4.
3.3 Loss function
We minimize a weighted sum of two loss functions: (1) eigenvector loss and (2) energy loss.
Let $\hat{U}$ denote a matrix of column vectors $\hat{u}_{i}$, where each $\hat{u}_{i}$ denotes the $i$th predicted eigenvector.
Eigenvector loss, used in [10], measures the difference between each $L\hat{u}_{i}$ and $\lambda_{i}\hat{u}_{i}$:
$\mathcal{L}_{eigvec}=\frac{1}{k}\lVert(L\hat{U}-\hat{U}\Lambda_{k})\rVert$
Energy loss, used in [8, 9, 11], aims to minimize the sum of Rayleigh quotients:
@@ 3.3 Loss function
To ensure the model does not output $k$ copies of the trivial eigenvector, we impose orthogonality on the final outputs of the model via QR decomposition, as proposed in [8].
We show that our loss functions satisfy necessary sign and basis invariances. Full proofs can be found in A.3.
3.4 Pre-Training Algorithm
Algorithm 1 Eigenvector Learning
@@ 3.4 Pre-Training Algorithm
Figure 1: Visual illustration of Laplacian Eigenvector GNN pre-training framework.
4 Experimental Results
4.1 Standalone pre-training
We pre-train a standard Graph Isomorphism Network (GIN) [12] on the eigenvector-learning task with the graph-level MLP head. Once the model has been pre-trained, we replace the graph-level MLP head with a downstream prediction MLP and fine-tune model weights. We evaluate our pre-training framework on two molecular datasets, ZINC-12k [13] and QM9 [14]. We record results of our experiments in Table 1. On 7 of 8 total targets, eigenvector pre-training substantially improves downstream performance.
ZINC QM9 Model Penalized $\log p$ $\mu$ $\alpha$ $\varepsilon_{\text{HOMO}}$ $\varepsilon_{\text{LUMO}}$ $\Delta_{\varepsilon}$ $\braket{R^{2}}$ ZPVE Pre-Trained GIN 0.353 0.484 0.489 0.00353 0.00371 0.00513 28.103 0.000477 Baseline GIN 0.438 0.472 1.132 0.00386 0.00399 0.00562 50.909 0.002400
4.2 Augmenting existing pre-training method
We augment the existing molecular pre-training methods proposed in [3] with eigenvector-learning. In particular, [3] proposes node-level pre-training tasks (context prediction and masking) on ZINC15 [13], followed by a graph-level supervised pre-training task on ChEMBL [15, 16]. We augment the graph-level supervised pre-training step by adding an additional MLP head to predict eigenvectors, and we evaluate on five downstream datasets based on [4].
Table 2: Test ROC-AUC (%, $\uparrow$) performance on 5 molecular prediction tasks when augmenting an existing pre-training method on a GIN base model. Sup. refers to the original supervised pre-training as implemented in [3], and Sup.+ refers to supervised training with additional eigenvector-learning (our method). Results for no pre-training are taken directly from [4]. All methods are tuned over seven learning rates and averaged over three seeds.
Dataset BACE BBBP Tox21 ToxCast SIDER Pretrain method MLP Head ContextPred, Sup.+ Graph-level $79.62\pm 3.63$ $\mathbf{70.76\pm 1.64}$ $77.94\pm 0.11$ $\mathbf{66.13\pm 0.34}$ $60.05\pm 0.99$ ContextPred, Sup.+ Node-wise $75.87\pm 3.11$ $68.74\pm 1.07$ $78.86\pm 0.06$ $63.78\pm 0.32$ $59.83\pm 0.53$ ContextPred, Sup. - $\mathbf{84.98\pm 1.28}$ $68.25\pm 0.48$ $77.44\pm 0.19$ $64.01\pm 0.81$ $\mathbf{62.87\pm 0.89}$ Masking, Sup.+ Graph-level $80.71\pm 3.84$ $68.33\pm 0.89$ $\mathit{79.09\pm 0.25}$ $\mathit{65.96\pm 0.20}$ $\mathit{62.41\pm 1.77}$ Masking, Sup.+ Node-wise $\mathit{81.02\pm 1.67}$ $\mathit{69.94\pm 1.76}$ $\mathbf{79.33\pm 0.41}$ $65.14\pm 0.44$ $59.38\pm 1.11$ Masking, Sup. - $75.42\pm 2.64$ $67.36\pm 4.60$ $78.33\pm 0.24$ $64.88\pm 0.82$ $61.6\pm 1.78$ No pre-training - $75.77\pm 4.29$ $69.62\pm 1.05$ $75.52\pm 0.67$ $63.67\pm 0.32$ $59.07\pm 1.13$
Table 3: Test MAE ($\downarrow$) performance on ZINC (12k subset) dataset when augmenting a graph structural encoder. All results using no structural encoder or the base GPSE are taken directly from [1]. GPSE+ refers to GPSE with modified eigenvector-learning (our method). Our experimental results are averaged over three seeds.
Base model Structural encoder MLP Head MAE GPS GPSE+ Graph-level $\mathbf{0.0629\pm 0.0016}$ GPSE+ Node-wise $0.0663\pm 0.0024$ GPSE - $\mathit{0.0648\pm 0.0030}$ None - $0.118\pm 0.005$ GIN GPSE+ Graph-level $0.1231\pm 0.0026$ GPSE+ Node-wise $0.1299\pm 0.0010$ GPSE - $0.124\pm 0.002$ None - $0.285\pm 0.004$
4.3 Augmenting existing graph structural encoder
We modify Graph Positional and Structural Encoder (GPSE) [1]. While GPSE already incorporates eigenvector-learning by including Laplacian positional encodings (LapPE) as a prediction target, GPSE uses MAE and cosine similarity loss on the absolute value of each eigenvector and a node-wise prediction head for every node property. We replace GPSE’s eigenvector-learning component with our own, using a separate MLP head and eigenvector loss. We keep all other GPSE model settings the same.
Following [1], we pre-train our modified GPSE model on MolPCBA [17]. We evaluate the effectiveness of these encodings by training both the Transformer model GPS [18] and a standard GIN, augmented with these encodings, on the downstream molecule property prediction task for the ZINC 12k [13] subset.
We report results on the effectiveness of these new learned encodings in Table 3. Modified eigenvector-learning with a graph-level MLP improves performance on downstream performance for both the GIN and GPS models, with our graph-level GPS+GPSE+ configuration achieving SOTA performance over all model and encoding configurations tested in [1]. We also demonstrate that our choice of loss function is crucial for eigenvector-learning in A.7.
5 Limitations and Future Work
There are several promising future directions toward improving the Laplacian eigenvector pre-training framework. We have demonstrated the effectiveness of the framework for pre-training and fine-tuning a GNN on the same dataset, or on domain-related datasets. However, we are yet to explore the effectiveness of eigenvector pre-training as a transfer learning framework.
Further, the practical implementation of the graph-level MLP requires adding padding to the concatenated node embeddings to accommodate for graphs of differing sizes. This creates an additional challenge for the MLP to learn meaningful relationships between the individual node embeddings within the graph-level vector. One could explore other ways of creating a rich graph-level representation while avoiding this challenge.
Appendix A Appendix
A.1 Full eigenvector pre-training pipeline
We provide a broad algorithmic outline of eigenvector pre-training process in Algorithm 2.
@@ A.1 Full eigenvector pre-training pipeline
A.2 Loss function

### v2

# A Graph Laplacian Eigenvector-based Pre-training Method for Graph Neural Networks

@@ Abstract
The development of self-supervised graph pre-training methods is a crucial ingredient in recent efforts to design robust graph foundation models (GFMs). Structure-based pre-training methods are under-explored yet crucial for downstream applications which rely on underlying graph structure. In addition, pre-training traditional message passing GNNs to capture global and regional structure is often challenging due to the risk of oversmoothing as network depth increases. We address these gaps by proposing the Laplacian Eigenvector Learning Module (LELM), a novel pre-training module for graph neural networks (GNNs) based on predicting the low-frequency eigenvectors of the graph Laplacian. Moreover, LELM introduces a novel architecture that overcomes oversmoothing, allowing the GNN model to learn long-range interdependencies. Empirically, we show that models pre-trained via our framework outperform baseline models on downstream molecular property prediction tasks.
@@ 1 Introduction
Graph Neural Networks (GNNs) have become a powerful tool in node and graph representation learning, with successful applications across domains ranging from biomedicine (Cantürk et al., 2023; Yan et al., 2024) to social networks (Fan et al., 2019). More recently, graph foundation models (GFMs) are emerging as an exciting field; inspired by the success of large language models (LLMs), researchers are exploring the possibility of creating large graph-based models with emergent capabilities across many domains (Liu et al., 2025; Xie et al., 2022; Wang et al., 2025).
A key ingredient in this effort is the creation of self-supervised tasks which can be performed on large unlabeled graph datasets (Liu et al., 2022). A variety of pre-training methods have been proposed, with the majority of such methods being feature-based, including contrastive losses and feature reconstruction (Liu et al., 2025; Xie et al., 2022). A few structure-based methods, which precompute labels based on graph topology, have been proposed (Peng et al., 2020a; Hwang et al., 2020). However, this category of pre-training approach, recently termed graph property prediction (Liu et al., 2025), remains under-explored. Furthermore, limitations of GNNs in capturing long-range structural information caused by oversmoothing and underreaching (Cai & Wang, 2020; Oono & Suzuki, 2019; Keriven, 2022) pose challenges when approaching graph property prediction tasks which rely on global structure.
We propose the Laplacian Eigenvector Learning Module (LELM) for GNNs and GFMs. The low-frequency Laplacian eigenvectors capture a range of global, regional, and local graph structure, making them well-suited as a structural target. LELM utilizes a novel global MLP prediction head during pre-training that allows the GNN model to learn long-range relationships without requiring excessively deep networks, and it augments pre-training data with positional features to overcome expressivity limits of GNNs. LELM is highly flexible: it can be used to pre-train any GNN to improve downstream performance across all kinds of graph-based datasets, and it can be used either as an independent pre-training method or as a plug-in addition to existing graph pre-training pipelines.
@@
    We propose LELM, a Laplacian eigenvector-based pre-training module for GNNs and GFMs.
@@
    We introduce a novel MLP head that enables GNNs to capture large-scale structure. Furthermore, we propose a set of augmented node features based on the graph diffusion operator.
@@
    We empirically demonstrate that LELM provides performance improvements over baseline models.
2 Background
Given a graph ${\mathcal{G}}=(V,E)$ with unnormalized adjacency matrix $\mathbf{A}$} and degree matrix $\mathbf{D}$}, the unnormalized Laplacian $\mathbf{L}$} of a graph $\mathcal{G}$} is defined as ${\mathbf{L}}={\mathbf{D}}-{\mathbf{A}}$.
Let $\lambda_{1},\lambda_{2},\dots,\lambda_{k}$ denote the $k$ lowest eigenvalues of $\mathbf{L}$} in nondecreasing order. Let $\psi_{1},\psi_{2},\dots\psi_{k}$ denote the corresponding eigenvectors, such that we have:
${\mathbf{L}}\mbox{$\mbox{$\psi$}$}_{i}=\lambda_{i}\mbox{$\mbox{$\psi$}$}_{i}$
The low-frequency graph Laplacian eigenvectors have been used for positional encodings, spectral GNNs, spectral clustering, and to generate provably minimal graph cuts. We provide examples and references for each of these applications in Section B.
@@ 3 Method
Figure 1: Overview of the LELM pre-training pipeline. Here, “Base GNN" and “Downstream Prediction Head" can be any user-defined model architecture.
The LELM pre-training framework consists of three primary components:
  • Node feature augmentation: We provide initial features based on the graph diffusion operator. These embeddings provide the base GNN model with additional structural context which is useful for the pre-training task as well as for downstream structure-aware tasks.
  • Graph-level MLP: We pass a graph-level aggregated representation from a base GNN into our prediction MLP head. The graph-level MLP overcomes the challenge of oversmoothing by allowing the base GNN to learn long-range node interactions without requiring an excessively deep message passing network.
  • Eigenvector prediction: During pre-training, we task the model to predict the $k$ lowest-frequency eigenvectors of the graph Laplacian.
Together, these components allow the base GNN to learn local and global graph structure while overcoming intrinsic expressivity limitations, making LELM a robust pre-training framework for structure-aware downstream tasks. A visual overview of our pipeline can be seen in Figure 1, and a detailed algorithmic overview of our pipeline can be found in Section C.
3.1 Node feature augmentation
We propose two kinds of embeddings: (1) wavelet positional embeddings, which encode relative positional information between nodes, and (2) diffused dirac embeddings, which encode local connectivity structures around each node. Both embeddings use the graph diffusion operator, and capture local aggregate information on each node. We provide more details as well as theoretical guarantees in Section G.
3.2 Graph-level MLP
Once we perform node augmentation, we pass the augmented graph as input into a base GNN model and graph-level MLP to output predicted eigenvectors.
@@ 3.2 Graph-level MLP
Graph-level MLP: We concatenate the node-wise output of the base GNN model to form a graph-level aggregated representation. We then pass the aggregated vector through an MLP model to produce the low-frequency Laplacian eigenvectors.
Eigenvector-learning methods for other applications (such as spectral clustering) use a node-wise MLP head, processing each node’s eigencoordinates independently based on their learned hidden embedding (Shaham et al., 2018; Dwivedi et al., 2021; Cantürk et al., 2023). Such methods fail to address oversmoothing as a node-wise MLP cannot learn long-range interactions; instead, one must use several layers of message passing within the base GNN (Cantürk et al., 2023).
3.3 Eigenvector prediction
During pre-training, the model output aims to minimize a weighted sum of two loss functions: (1) eigenvector loss and (2) energy loss.
To ensure the model does not output $k$ copies of the trivial eigenvector, we impose orthogonality on the final outputs of the model via QR decomposition, as proposed by Shaham et al. (2018). Let each $\hat{{{\mathbf{u}}}}_{i}$ denote the $i$th predicted eigenvector via LELM after orthogonality has been imposed.
Energy loss, used in Shaham et al. (2018); Dwivedi et al. (2021); Ma & Zhan (2023), computes the sum of Rayleigh quotients:
$\mathcal{L}_{energy}=\frac{1}{k}\sum_{i=1}^{k}\hat{{{\mathbf{u}}}}_{i}^{\top}{\mathbf{L}}\hat{{{\mathbf{u}}}}_{i}$
Energy loss is minimized when the predicted eigenvectors span the same subspace as the ground-truth eigenvectors; to enforce a strict ordering and basis on the predicted eigenvectors, we additionally impose eigenvector loss, used in Mishne et al. (2019):
$\mathcal{L}_{eigvec}=\frac{1}{k}\sum_{i=1}^{k}\lVert{\mathbf{L}}\hat{{{\mathbf{u}}}}_{i}-\lambda_{i}\hat{{{\mathbf{u}}}}_{i}\rVert$
We discuss further intuition behind our loss functions in Section D and provide proofs for necessary sign and basis invariances in Section F.
4 Experimental Results
To evaluate the effectiveness of LELM, we conduct various experiments on real-world datasets.
Comparison against baseline models: We pre-train a standard Graph Isomorphism Network (GIN) (Xu et al., 2019) and GPS (Rampášek et al., 2022), a graph transformer, using LELM. Once the model has been pre-trained, we replace the graph-level MLP head with a downstream prediction MLP and fine-tune model weights. We evaluate our pre-training framework on three molecular datasets ZINC, ZINC-12k (Sterling & Irwin, 2015) and QM9 (Ramakrishnan et al., 2014). For each of these models, we compare LELM against the same GNN model without pre-training. For both models, pre-training improves performance for all but one of the downstream targets. We record results of our experiments in Table 1.
@@ 4 Experimental Results
Comparison against alternative pre-training targets: We compare LELM to alternative structure-based pre-training targets, including node degree, local clustering coefficient, cycle counting, and Laplacian eigenvalues. For each alternative target, we pre-train a standard GIN (with our default settings) using L1 loss. Results are recorded in Table 2. Here, we see that LELM outperforms alternative structure-based pre-training targets, indicating that the Laplacian eigenvectors are a better pre-training target than many other natural choices for graph targets.
Table 2: Test MAE ($\downarrow$) performance comparison when pre-training a GIN on ZINC to predict alternative structural training targets.
[…]
Ablation on MLP head: We pre-train both the GIN and GPS using LELM, replacing the graph-level MLP with a standard MLP which acts on each node’s embeddings individually. Results can be seen in Table 3; for both the GIN and GPS, pre-training is more effective with the graph-level MLP.
Table 3: Test MAE ($\downarrow$) performance comparison on ZINC when replacing the graph-level MLP with a basic node-wise MLP during pre-training.
[…]
Enhancing existing pre-training methods: To demonstrate the flexibility of our method, we complement an existing pre-training method with LELM and demonstrate performance improvements in the majority of downstream tasks. Full experimental details and results can be found in Section I.
5 Limitations and Future Work
There are several promising future directions toward improving the LELM pre-training framework. First, we have demonstrated the effectiveness of the framework for pre-training and fine-tuning a GNN on the same dataset, or on domain-related datasets, but we have yet to explore the effectiveness of LELM as a transfer learning framework. A future direction could be exploring the viability of LELM when pre-training on synthetic graphs or on cross-domain datasets.
Appendix A Related works: GNN pre-training methods
Towards the goal of improving graph foundation models, a variety of self-supervised graph pre-training tasks have been proposed. According to the taxonomy provided by Liu et al. (2025); Xie et al. (2022), existing graph pre-training methods can be categorized into two broad categories: contrastive and predictive methods.
Contrastive methods maximize mutual information between pairs of data views using objectives like Jensen-Shannon estimator (Nowozin et al., 2016) or InfoNCE (Oord et al., 2018). Methods can be categorized by the types of views used: graph-node (Sun et al., ; Veličković et al., ; Peng et al., 2020b), subgraph-node (Hu et al., ; Jiao et al., 2020), and subgraph-subgraph Qiu et al. (2020). Some methods also employ graph augmentation to generate two views (You et al., 2020).
Predictive methods, also referred to as generative methods (Liu et al., 2025), self-generate labels and train to predict these labels. A first class of predictive models uses graph reconstruction, whether by using node/edge masking (Xie et al., 2020; Batson & Royer, 2019; Hu et al., ) or using autoencoders (Wang et al., 2017; Kipf & Welling, 2016). A second class of predictive methods are property prediction methods, which precompute underlying graph properties as labels. Examples include statistical properties such as k-hop connectivity (Peng et al., 2020a) or topological properties like a meta-path (Hwang et al., 2020). Overall, there are a lack of works on property prediction-based methods, with the majority of predictive pre-training methods falling under the former category of graph reconstruction (Liu et al., 2025). Our method, LELM, is the first property prediction method to use the graph Laplacian eigenvectors as a pre-training target.
Appendix B Graph Laplacian eigenvector applications
Provably minimal graph cuts: The second-lowest eigenvector $\psi_{2}$, known as the Fiedler vector, can be used to generate a provably “good" cut on a graph; in particular, for some arbitrary threshold $s\in\mathbb{R}$, we can define a Fiedler cut $C$ to be:
$C=(\{i:\psi_{2}(i)<s\},\{i:\psi_{2}(i)\geq s\})$
On any bounded-degree $n$-vertex planar graph, the optimal Fiedler cut has ratio $O(\frac{1}{n})$ (Spielman & Teng, 1996).
Positional encodings: The low-frequency Laplacian eigenvectors naturally encode a global position on the graph. As a result, Laplacian positional encodings (LapPE) have been used as a standard positional encoding for graph transformer models (Dwivedi & Bresson, 2020; Rampášek et al., 2022). In practice, directly using the Laplacian eigenvectors as positional encodings creates sign and basis ambiguity issues, as $\psi_{i}$ is an eigenvector of $L\iff-\psi_{i}$ is an eigenvector of $L$. Approaches to solving this problem include designing an architecture which processes the Laplacian eigenvectors in a sign- and basis-invariant manner (Lim et al., 2022) or defining canonical directions for the eigenvectors (Ma et al., 2023).
Spectral GNNs: A variety of methods, known as spectral graph neural networks, use the Laplacian eigendecomposition to learn filters in signal domain (Bo et al., 2023b). Some methods explicitly compute or approximate the $k$ lowest-frequency Laplacian eigenvectors, learning advanced filters on the corresponding eigenvalues (Bruna et al., 2013; Liao et al., 2019; Bo et al., 2023a). Other methods instead learn polynomial filters on the graph (Defferrard et al., 2016; He et al., 2022), circumventing the expensive process of eigendecomposition by learning a $k$-degree polynomial function with respect to $L$, i.e. $p(L)=\theta_{0}I+\theta_{1}L+\dots+\theta_{k}L^{k}$.
Spectral clustering: The Laplacian eigenvectors have also been used for clustering applications. Given a set of data $x_{1},\dots x_{n}$, Belkin & Niyogi (2001) construct a weighted graph $G$ with $n$ nodes using a heat kernel. Then to generate a $k$-dimensional embedding, Belkin & Niyogi (2001) compute the $k$-lowest eigenvectors $\psi_{2},\dots\psi_{k+1}$ (omitting the trivial eigenvector) of the graph Laplacian and assign data point $x_{i}$ the embedding $(\psi_{2}(i),\psi_{3}(i),\dots,\psi_{k+1}(i))$. Shaham et al. (2018); Chen et al. (2022) learn this spectral map using a neural network, allowing for a natural extension of this map to new datapoints.
Appendix C Complete algorithms
We provide a detailed overview of the eigenvector pre-training process in Algorithm 1, and an overview of the full pre-training and fine-tuning pipeline in Algorithm 2.
Algorithm 1 Eigenvector Prediction
@@ Appendix C Complete algorithms
Appendix D Loss function
We minimize a weighted sum of two loss functions: (1) eigenvector loss and (2) energy loss. Both loss functions respect necessary sign and basis invariances of Laplacian eigenvectors; full proofs can be found in F.
Energy loss, used by Shaham et al. (2018); Dwivedi et al. (2021); Ma & Zhan (2023), aims to minimize the sum of Rayleigh quotients:
[…]
This loss function is motivated by the iterative optimization problem following from Courant-Fischer, which states that the eigenvectors of $L$ (and the eigenvectors of any Hermitian matrix) can be equivalently expressed as solutions to the following iterative optimization problem:
$\psi_{k}\in\mathop{\operator@font arg\,min}_{\begin{subarray}{c}\|x\|=1\\ x\perp\psi_{1},\dots,\psi_{k-1}\end{subarray}}x^{\top}Lx.$
However, minimizing this loss function only minimizes the sum of the first $k$ Rayleigh quotients, meaning the minimizer (subject to orthogonality) is any set of vectors spanning same subspace spanned by the $k$ lowest frequency eigenvectors. For applications in clustering, this is reasonable, as the exact basis in which embeddings are expressed is often irrelevant; however, to require the model to truly predict the $k$-lowest eigenvectors, we must include a more explicit penalty, such as eigenvector loss.
Eigenvector loss, used by Mishne et al. (2019), measures the difference between each $L\hat{u}_{i}$ and $\lambda_{i}\hat{u}_{i}$:
$\mathcal{L}_{eigvec}=\frac{1}{k}\lVert(L\hat{U}-\hat{U}\Lambda_{k})\rVert$
Eigenvector loss enforces both the correct basis and a strict ordering (up to eigenvalue multiplicity) on the predicted eigenvectors. Our final loss function is then:
$\mathcal{L}=\alpha\cdot\mathcal{L}_{energy}+\beta\cdot\mathcal{L}_{eigvec}$
Appendix E Loss function alternative expressions